X Y Z a b cin Cout Cout Find the minimal SOP and POS representation of this circuit.Implement the circuit from the previous question using only a decoder of your choice in addition to one extra gate of your ch
Answer #1

= 6 x 6 z coot = X Y ty z + ZX HA 5 = ab НА cout = ab where a= st z 6 cout b= y cout F = x+st cout HA =xt a obtab = x+ ab + a

ху я Ч4+ 4+ + + + + + + + (4+4+) о - (3 7 6 (14+7 +ъ) - 7( xy +2+) +4 (14+чата) з - за 4 ч. за ..) - уз 2 ( 3) (+3). 13) - 1

+ 2 F = x+ łý z + 2y Z + x 9 žtay Etayz 11 @ y. cout = y. By+ y2 + 2x) = xy + yz txyz ③ = xy, Jz. xyz = (x+5) (+2) (a +3 +Z)F = Z at 9 of n 15 1 1 o ý yz Lazy L 1 1 1 az F = Em (0, 1, 2, 4, 5, 6, 7) 3x8 de coda By la A using a 3x8 Decodue Enable